OpenStudy (anonymous):

-4p-10(2p-3)=22-7(1-2p)

OpenStudy (anonymous):

p= 15/38

OpenStudy (anonymous):

so can you work the whole problem out

OpenStudy (anonymous):

1) -4p - 10(2p - 3) = 22 -7 (1-2p) 2) -4p - 20p + 30 = 22 - 7 + 14p 3) -24p + 30 = 15 + 14p 4) 15 = 38p 5) p = 15/38